# When Collaborative Maintenance Falls Short: The Persistence of Retracted Papers on Wikipedia

### v2

@@ 2.3 Retractions, Online Attention, and Challenges of Correction
Retractions are intended to correct the scientific record, but their corrective effects are often limited once flawed research has circulated online. Prior work shows that retracted papers frequently receive more online attention compared to their non-retracted counterparts, particularly on social media, news platforms, and in knowledge repositories peng2022dynamics . This attention is often exhausted before the retraction is issued, and, in some cases, the retraction itself may unintentionally amplify visibility. While increased attention can surface public criticism—especially on platforms like Twitter/X peng2022dynamics ; amiri2024tweeters —only a small and selective group of users tends to engage in such critical discussions. Moreover, retracted papers have also been shown to disseminate false information, particularly around sensitive topics such as vaccination, thereby undermining public trust in science abhari2024they . Emerging work has explored whether language models can help flag problematic papers based on these online signals, with promising results using Twitter/X data zheng2024can .
@@ 3.1 Data Collection
Figure 2: Overview of the data collection and annotation process. The left panel shows the construction of the main dataset, including the extraction of 1,181 retracted paper–Wikipedia citation pairs from Crossref Event Data and corresponding revision histories from Wikipedia. Manual annotation determined whether each citation was cited with its retraction status noted at its first appearance and corrected in the latest revision. The right panel displays the separate dataset of all Wikipedia-cited papers (N = 1,590,982), which is used to calculate citation ratios across academic domains. Retraction status is linked using Retraction Watch, and domains are categorized via OpenAlex.
@@
We integrated data from Retraction Watch oransky2012retraction , Crossref Events hendricks2020crossref , Altmetric LLC altmetric , OpenAlex priem2022openalex , as well as Wikipedia’s page view statistics (Wikimedia REST API) and revision histories (MediaWiki Action API). This integration enables us to systematically examine how retracted research is cited and handled on Wikipedia, allowing us to investigate further how features of papers, Wikipedia pages, online mentions, and temporal trends (e.g., whether citations occurred before or after retraction, and the time between publication and retraction) are associated with the correction of retracted paper citations.
@@
The Retraction Watch database is the most comprehensive and widely adopted resource on retracted articles brainard2018rethinking , indexing all retracted Digital Object Identifiers (DOIs) and their metadata, such as retraction date, publication date, and retraction reason. In this study, we used Retraction Watch to obtain a list of retracted papers and their associated metadata, including article titles, retraction dates, publication dates, and retraction reasons. To ensure our analysis focused exclusively on retracted papers, we excluded the ones with expressions of concern or corrections. Expressions of concern serve as warnings while investigations are ongoing, and corrections address minor issues without invalidating the entire study. Retractions, by contrast, formally withdraw a paper from the academic record.
@@
We used the Crossref Event API to collect all Wikipedia pages that have cited the retracted DOIs. Crossref periodically crawls Wikipedia and logs pages that contain outbound links to Crossref DOIs, along with detection timestamps and metadata hendricks2020crossref . The Crossref Event API is the most accurate in indexing Wikipedia citation data ortega2018reliability , and only indexes papers with valid DOIs. 833 retracted papers were cited on 900 Wikipedia pages, forming 1,181 retracted paper–Wikipedia page citation pairs. To accurately date when the retracted DOI was introduced to the Wikipedia pages and when the retracted DOI was properly corrected on Wikipedia (complete removal or adding an in-text warning), we used the MediaWiki Action API to retrieve and analyse all revisions ($N=824,012$) of these pages. This allowed us to perform a detailed temporal analysis of how retracted papers were cited and subsequently addressed on Wikipedia. To ensure our analysis was relevant to public-facing content, we excluded pages from non-article namespaces such as Talk, Sandbox, Draft, and User.
@@
We constructed a dataset separate from the one used in our main analysis to compare the academic domain distribution of retracted and non-retracted papers cited on Wikipedia. For consistency, we identified 538 retracted papers using the latest Altmetric data dump (current as of August 14, 2023), rather than relying on our original Wikipedia retracted citations dataset. This allowed us to align data sources when comparing them to the 1,590,444 non-retracted papers also cited on Wikipedia and indexed in Altmetric. The Crossref Event API proved impractical for extracting such a large volume of citations to non-retracted papers, so relying on Altmetric for both retracted and non-retracted papers ensured source consistency for this descriptive analysis. This dataset is used solely for comparing academic domain distributions and is not part of our modeling of correction outcomes. To enrich the dataset, we retrieved academic domain classifications from OpenAlex, along with additional metadata such as open access status, citation counts, and co-author team sizes. All data were collected in October 2024.
@@ 3.2 Types of Retracted Paper Citations
  • Cited After Retraction With In-Text Acknowledgment. In some cases, retracted papers are introduced with explicit acknowledgment of their retraction status. This definition intentionally excludes automated warnings in the reference section, as our focus is on the human editorial actions that make a retraction’s status transparent to a reader of the main article text.
@@ 3.3 Handling Problematic Retracted Paper Citations
By manually inspecting the latest version of pages containing problematic citations in our dataset, we observe three ways of handling problematic retracted paper citations on Wikipedia:
  • Complete Removal. The most straightforward approach to correcting citations of retracted papers has been to remove them entirely from the Wikipedia article. This method eliminates any potential for future misinformation stemming from the problematic source. However, it can be controversial if the retracted paper contributed historically significant or contextually relevant information to the article. In academic paper writing, the complete removal of retracted paper citations is always chosen when editors and reviewers do not consider its continued presence essential retractionwatch2018how .
  • In-Text Retraction Acknowledgment. A second approach involves placing a clear warning near the citation in the Wikipedia article text, ideally in the same paragraph where the retracted paper is referenced. This method informs readers about the paper’s retracted status while preserving the citation for transparency or historical context. Such warnings may appear as editorial notes or parenthetical remarks, e.g., “(This paper has been retracted).” This category also includes other explicit in-text editorial flags, such as the better source needed template. This strategy balances accuracy and informational completeness, but it relies heavily on editors’ consistency and awareness.
  • Reference Section Acknowledgment Only. In many cases, a paper’s retracted status is only acknowledged in the reference section, without modifying the article’s main text. Because these reference-level warnings are less visible to casual readers—who may not consult the references—this practice allows potentially invalidated claims to persist unchallenged in the main body of the article. As a result, critical misinformation can remain embedded in Wikipedia’s content. For example, automated tools like RetractionBot Headbomb2024Signpost are designed to flag retracted papers on Wikipedia by adding annotations primarily to the reference section.
For our analysis, we consider citations to be corrected if it was Complete Removal or In-Text Retraction Acknowledgment, as these approaches offer high visibility to readers.
To identify these two types of corrections, we employed a two-step process. First, to find instances of Complete Removal, we searched the revision history of each article for the titles and DOIs of the retracted papers. Second, for all remaining cases where the citation was still present, we manually reviewed the most recent revision to identify instances of In-Text Retraction Acknowledgment, assessing whether the retraction status was explicitly mentioned in the main text.
@@ 3.4 Analytical Approach: Survival Analysis
Our analysis seeks to understand the factors that influence how quickly problematic citations are corrected. A significant portion of these citations remained uncorrected at the end of our data collection period. This type of incomplete observation is known as right-censored data. Standard regression models are not suited to handle such data. Therefore, to properly account for these censored observations, we employ survival analysis, a method well-suited for modeling time-to-event data in online communities ma2017write ; mindel2025timely . This method enables us to model the time until a correction occurs (the “event”) while appropriately incorporating information from both corrected and uncorrected (censored) citations.
@@ 3.5 Dependent Variable: Time to Correction
Crucially, the start of the risk period for each citation is defined as the later of two dates: the paper’s official retraction date or the date the citation was first added to the Wikipedia page. This ensures that a citation is only considered “at risk” after it is both present on Wikipedia and officially retracted. As a data cleaning step, we excluded a small number of “proactive corrections”—cases where a correction was made before the official retraction date. To avoid numerical issues in parametric survival fitting, all durations were adjusted by adding a small $\epsilon=10^{-3}$ days.
A correction event is defined as either the complete and persistent removal of the citation or the addition of an in-text warning of the retraction in the immediate context of the citation. For citations that remained uncorrected as of our data collection endpoint (October 24, 2024), they are treated as right-censored observations. Their duration is the time from the start of their risk period until this endpoint.
@@ 3.5.1 Correction via In-Text Retraction Acknowledgment
For citations that were eventually corrected by mentioning the retraction status in the main text of the article, we begin by identifying a retraction-related keyword in the final, corrected revision of the Wikipedia article. These may include general terms like “retracted” or “withdrawn,” or more specific phrases such as “The article was retracted on 13 August 2010.” Using this final revision, we select a representative keyword or phrase associated with the correction and trace backward through the revision history to identify the earliest revision where that keyword appeared.
To avoid mistaking short-term rewording or editing noise as correction reversals, we required that the identified correction keyword or phrase, once added, remain in subsequent revisions, allowing for a temporary disappearance of up to three consecutive revisions. This heuristic balances robustness against transient noise with sensitivity to meaningful editorial behavior. Because articles may cite multiple retracted papers, we manually verify that the keyword refers to the specific citation of interest. If needed, we extract a more precise phrase from the later revision that unambiguously refers to the retracted paper in question.
@@ 3.5.3 Assumptions
  • Once corrected, a citation remains corrected in that same form throughout subsequent revisions.
  • The retraction-related keyword or phrase, once added, remains consistent in subsequent revisions.
@@ 3.6 Independent Variables
Table 1: Summary statistics of independent variables used in the survival analysis. Variables marked with “(at risk start)” were measured at the beginning of each citation’s unique risk period, which starts at the later of two dates: the paper’s official retraction date or the date the citation was added to Wikipedia.
@@
   Open Access 0.62 0.48 0 1 1
   Academic Citation Count (at risk start) 147.77 417.75 0 5436 46.00
   Team Size 6.34 4.65 0 28 5.00
@@
   Tweet Mention Count (at risk start) 345.26 2159.96 0 33946 2.00
@@
   Time to Retraction (days) 2409.45 2096.41 0 11344 1914.50
@@
   Retraction Reason (salient) 0.68 0.74 0 2 1.00
@@
   Number of Unique Editors (at risk start) 151.79 275.27 0 2343 34.50
@@
To identify factors that accelerate or decelerate the correction of retracted paper citations on Wikipedia, we examine four different groups of characteristics, as follows (see Table 1):
@@ 3.6.1 Paper Features
Paper characteristics are particularly important because they capture the intrinsic properties of the publications that may either facilitate or hinder their detection and correction. These data were retrieved from external scholarly databases, including OpenAlex priem2022openalex , Retraction Watch oransky2012retraction , and Altmetric altmetric .
@@
  • Open Access. Open access status is coded as a binary variable. This variable allows us to examine whether open-access availability affects the speed with which retracted papers are corrected on Wikipedia. Prior research suggests that open-access papers may receive greater visibility and more citations due to their unrestricted availability piwowar2018state ; teplitskiy2017amplifying . Zheng et al. found that open access may help speed up the detection and retraction of flawed articles in biochemistry zheng2024can .
@@
  • Academic Citation Count. Academic citations serve as a proxy for a paper’s scholarly impact and visibility. To avoid look-ahead bias, we measured this variable dynamically for each citation pair. Specifically, we only counted citations the paper had accumulated before the start of its risk period on Wikipedia, ensuring our measure of scholarly impact could not be influenced by the correction process itself. Using the OpenAlex API, we determined the number of citations a retracted paper had received up to the start of each citation’s unique risk period (i.e., the point after it was both retracted and cited on Wikipedia). While a paper’s higher citation count could increase its visibility and lead to a faster correction on Wikipedia van2011science ; wang2021science , it could also create editorial inertia if its findings are widely accepted, resulting in a slower correction.
@@
  • Team Size. Team size is measured by counting the number of paper authors. Larger teams may contribute to greater research visibility and faster dissemination due to broader professional networks li2022untangling . However, prior studies suggest that increased team size can also complicate accountability, potentially leading to delays in retraction or correction processes mongeon2016costly .
@@
  • Academic Domains. We include academic domains to examine whether the speed of citation correction varies across broad fields of research. Domains are derived from OpenAlex’s primary_topic field, which is based on Scopus’s ASJC classification Elsevier and assigned using OpenAlex’s topic classification methods openalexend . We focus on four high-level domains: Health Sciences, Life Sciences, Physical Sciences, and Social Sciences. To incorporate these categorical variables into our regression model, we applied one-hot encoding and omitted Health Sciences as the reference category to avoid multicollinearity. Differences across domains may reflect disparities in editorial practices, public scrutiny, or disciplinary norms. For example, Health Sciences often receive greater media and public attention, which can accelerate the detection and correction of retracted work furman2012governing . In contrast, corrections in the Social Sciences may lag due to longer publication cycles.
@@
  • Tweet Mention Count. Tweet mention count data is collected from Altmetric LLC altmetric . To ensure methodological consistency and avoid look-ahead bias, we measured the number of tweets up to the start of each citation’s unique “at risk” period. This count serves as a proxy for the paper’s online visibility and public exposure at the moment it became problematic on Wikipedia. While a higher count of these pre-existing tweets could signal a more prominent paper that editors are more likely to scrutinize, we acknowledge that the attention generated by the retraction event itself is another important mechanism. Although modeling this second effect is a valuable direction for future work, this study focuses on the role of a paper’s pre-existing visibility. Prior research highlights the role of online engagement in scientific controversies, finding that public attention on platforms like Twitter often spikes around the time a paper is retracted peng2022dynamics . While news and blog mentions also serve as valid proxies for online engagement, we found they were highly correlated with the Tweet mention count. To avoid multicollinearity in our model, we therefore selected the Tweet mention count as the most representative measure.

### v3

@@ 2.3 Retractions, Online Attention, and Challenges of Correction
Retractions are intended to correct the scientific record, but their corrective effects are often limited once flawed research has circulated online. Prior work shows that retracted papers frequently receive more online attention compared to their non-retracted counterparts, particularly on social media, news platforms, and in knowledge repositories Peng et al. (2022). This attention is often exhausted before the retraction is issued, and, in some cases, the retraction itself may unintentionally amplify visibility. While increased attention can surface public criticism—especially on platforms like Twitter/X Peng et al. (2022); Amiri et al. (2024)—only a small and selective group of users tends to engage in such critical discussions. Moreover, retracted papers have also been shown to disseminate false information, particularly around sensitive topics such as vaccination, thereby undermining public trust in science Abhari and Horvát (2024). Emerging work has explored whether language models can help flag problematic papers based on these online signals, with promising results using Twitter/X data Zheng et al. (2025).
@@ 3.1 Data Collection
Figure 2: Overview of the data collection and annotation process. The left panel shows the construction of the main dataset, including the extraction of 1,181 retracted paper–Wikipedia citation pairs from Crossref Event Data and corresponding revision histories from Wikipedia. One of the authors manually annotated whether each citation was cited with its retraction status noted at its first appearance and corrected in the latest revision (removal or retraction status noted on the page). The right panel displays the separate dataset of all Wikipedia-cited papers (N = 1,590,982), which is used to calculate citation ratios across academic domains. Retraction status is linked using Retraction Watch, and domains are categorized via OpenAlex.
@@
We integrated data from Retraction Watch Oransky and Marcus (2012), Crossref Events Hendricks et al. (2020), Altmetric LLC Altmetric Support (2020), OpenAlex Priem et al. (2022). We also used the Wikimedia REST API to retrieve monthly page view statistics for articles citing retracted papers, and the MediaWiki Action API to analyze their full revision histories. This integration enables us to systematically examine how retracted research is cited and handled on Wikipedia, allowing us to investigate further how features of papers, Wikipedia pages, online mentions, and temporal trends (e.g., whether citations occurred before or after retraction, and the time between publication and retraction) are associated with the correction of retracted paper citations.
@@
The Retraction Watch database is the most comprehensive and widely adopted resource on retracted articles Brainard (2018), indexing all retracted Digital Object Identifiers (DOIs) and their metadata, such as retraction date, publication date, and retraction reason. In this study, we used Retraction Watch to obtain a list of retracted papers and their associated metadata, including article titles, retraction dates, publication dates, and retraction reasons. To ensure our analysis focused exclusively on retracted papers, we excluded the ones with expressions of concern or corrections. Expressions of concern serve as warnings while investigations are ongoing, and corrections address minor issues without invalidating the entire study. Retractions, by contrast, formally withdraw a paper from the academic record.
@@
We used the Crossref Event API to collect all Wikipedia pages that have cited the retracted DOIs. Crossref periodically crawls Wikipedia and logs pages that contain outbound links to Crossref DOIs, along with detection timestamps and metadata Hendricks et al. (2020). The Crossref Event API is the most reliable in indexing Wikipedia citation data Ortega (2018), and only indexes papers with valid DOIs. 833 retracted papers were cited on 900 Wikipedia pages, forming 1,181 retracted paper–Wikipedia page citation pairs. To accurately date when the retracted DOI was introduced to the Wikipedia pages and when the retracted DOI was properly corrected on Wikipedia (complete removal or adding an in-text warning), we used the MediaWiki Action API to retrieve and analyse all revisions ($N=824,079$) of these pages. This allowed us to perform a detailed temporal analysis of how retracted papers were cited and subsequently addressed on Wikipedia. To ensure our analysis was relevant to public-facing content, we excluded pages from non-article namespaces such as Talk, Sandbox, Draft, and User.
@@
We constructed a dataset separate from the one used in our main analysis to compare the academic domain distribution of retracted and non-retracted papers cited on Wikipedia. For consistency, we identified 538 retracted papers using the latest Altmetric data dump (current as of August 14, 2023), rather than relying on our original Wikipedia retracted citations dataset. This allowed us to align data sources when comparing them to the 1,590,982 non-retracted papers also cited on Wikipedia and indexed in Altmetric. The Crossref Event API proved impractical for extracting such a large volume of citations to non-retracted papers, so relying on Altmetric for both retracted and non-retracted papers ensured source consistency for this descriptive analysis. This dataset is used solely for comparing academic domain distributions and is not part of our modeling of correction outcomes. To enrich the dataset, we retrieved academic domain classifications from OpenAlex, along with additional metadata such as open access status, citation counts, and co-author team sizes. All data were collected in October 2024.
@@ 3.2 Types of Retracted Paper Citations
  • Cited After Retraction With In-Text Acknowledgment. In some cases, retracted papers are introduced with explicit acknowledgment of their retraction status. This definition intentionally excludes automated warnings in the reference section, as our focus is on the human editorial actions that make a retraction’s status transparent to a reader of the main article text. At the same time, we treat RetractionBot flagging as an observable feature that may be associated with subsequent reader-facing repair.
@@ 3.3 Handling Problematic Retracted Paper Citations
By manually inspecting the latest version of pages containing problematic citations in our dataset, we distinguish three reader-relevant ways that retraction status may be handled in an article:
  • Complete Removal. The most straightforward approach to correcting citations of retracted papers has been to remove them entirely from the Wikipedia article. This method eliminates any potential for future misinformation stemming from the problematic source. However, it can be controversial if the retracted paper contributed historically significant or contextually relevant information to the article. In academic paper writing, the complete removal of retracted paper citations is often preferred when editors and reviewers do not consider its continued presence essential Watch (2018).
  • In-Text Retraction Acknowledgment. A second approach involves placing a clear warning near the citation in the Wikipedia article text, ideally in the same paragraph where the retracted paper is referenced. This method informs readers about the paper’s retracted status while preserving the citation for transparency or historical context. Such warnings may appear as editorial notes or parenthetical remarks, e.g., “(This paper has been retracted).” This category also includes other explicit in-text editorial flags, such as the better source needed template. We treat such editorial flags as a form of reader-facing repair because they explicitly signal sourcing problems in the main text, even if they do not name the retraction. This strategy balances accuracy and informational completeness, but it relies heavily on editors’ consistency and awareness.
  • Reference Section Acknowledgment Only. In many cases, a paper’s retracted status is only acknowledged in the reference section, without modifying the article’s main text. Because these reference-level warnings are less visible to casual readers—who may not consult the references—this practice leaves the main text without a visible retraction signal, which may increase the risk that readers interpret the cited claim as uncontested. As a result, potentially outdated sourcing can remain in the main text without a reader-facing warning in Wikipedia’s content. For example, automated tools like RetractionBot Headbomb (2024) are designed to flag retracted papers on Wikipedia by adding annotations primarily to the reference section.
For our analysis, we consider citations to be corrected if they fall under Complete Removal or In-Text Retraction Acknowledgment, as these approaches offer high visibility to readers.
To identify these two types of corrections, we employed a two-step process. First, to find instances of Complete Removal, we searched the revision history of each article for the titles and DOIs of the retracted papers. We normalized DOI strings (e.g., URL prefixes and case) and used case-insensitive matching in wikitext to identify first appearance and subsequent absence. Second, for all remaining cases where the citation was still present, we manually reviewed the most recent revision to identify instances of In-Text Retraction Acknowledgment, assessing whether the retraction status was explicitly mentioned in the main text.
@@ 3.4 Analytical Approach: Survival Analysis
Our analysis examines which observable features are associated with the timing of reader-facing corrections for problematic citations. A significant portion of these citations remained uncorrected at the end of our data collection period. This type of incomplete observation is known as right-censored data. Standard regression models are not suited to handle such data. Therefore, to properly account for these censored observations, we employ survival analysis, a method commonly used to model time-to-event data in online communities Ma et al. (2017); Mindel et al. (2025). This method enables us to model the time until a correction occurs (the “event”) while appropriately incorporating information from both corrected and uncorrected (censored) citations.
@@ 3.5 Dependent Variable: Time to Correction
Crucially, the start of the risk period for each citation is defined as the later of two dates: the paper’s official retraction date or the date the citation was first added to the Wikipedia page. This ensures that a citation is only considered “at risk” after it is both present on Wikipedia and officially retracted. As a data cleaning step, we excluded a small number of “proactive corrections”—cases where a correction was made before the official retraction date. To avoid undefined values when durations are zero (e.g., when the citation and retraction occur on the same date) and for models that include $\log(t)$-based terms, all durations were adjusted by adding a small $\epsilon=10^{-3}$ days.
A correction event is defined as either (i) removal of the citation that is sustained under our three-revision tolerance rule, or (ii) the addition of an in-text retraction warning in the immediate context of the citation. For citations that remained uncorrected as of our data collection endpoint (October 24, 2024), they are treated as right-censored observations. Their duration is the time from the start of their risk period until this endpoint.
@@ 3.5.1 Correction via In-Text Retraction Acknowledgment
For citations that were eventually corrected by mentioning the retraction status in the main text of the article, we begin by identifying a retraction-related keyword in the final, corrected revision of the Wikipedia article. These may include general terms like “retracted” or “withdrawn,” or more specific phrases such as “The article was retracted on 13 August 2010.” Using this final revision, we select a representative keyword or phrase associated with the correction and trace backward through the revision history to identify the earliest revision where that keyword appeared. We select this phrase on a per-citation basis to match the local wording used for the correction, rather than relying on a single global keyword list.
To avoid mistaking short-term rewording or editing noise as correction reversals, we required that the identified correction keyword or phrase, once added, remain in subsequent revisions, allowing for a temporary disappearance of up to three consecutive revisions. We acknowledge that acknowledgments may be reworded or restructured across revisions; we discuss this measurement limitation and our validation-based corrections in Section 4.2.2 and Section 3.5.3. This heuristic balances robustness against transient noise with sensitivity to meaningful editorial behavior. Because articles may cite multiple retracted papers, we manually verify that the keyword refers to the specific citation of interest. If needed, we extract a more precise phrase from the later revision that unambiguously refers to the retracted paper in question.
@@ 3.5.3 Assumptions
  • Once corrected, a citation typically remains corrected in that same form throughout subsequent revisions, allowing for brief reversals consistent with our three-revision tolerance.
  • The retraction-related keyword or phrase, once added, is typically consistent in subsequent revisions, though it may be reworded or replaced (e.g., synonym substitution or text restructuring). In our manual validation (Section 4.2.2), we found that remaining mismatches primarily stem from such rewriting dynamics or shifts between In-Text Acknowledgment and Complete Removal; we corrected the identified cases in the final analytical dataset.
@@
3.5.4 Validation of correction timestamps.
Because correction timestamps are automatically derived from revision histories using the heuristics above, we manually validated a subset of cases and corrected identified mismatches caused by editorial rewriting or strategy switches (e.g., acknowledgment vs. removal). We report the validation design, sample sizes, and accuracy rates in Section 4.2.2. This validation targets the automatically inferred correction timestamps (from revision heuristics), rather than the separate manual classification of citation types.
@@ 3.6 Independent Variables
   Open Access 0.62 0.49 0 1 1
   Academic Citation Count (at risk start) 148.90 419.99 0 5444 46.00
   Team Size 6.35 4.65 0 28 5.00
@@
   Tweet Mention Count (at risk start) 346.15 2162.73 0 33946 1.50
@@
   Time to Retraction (days) 2414.00 2097.01 0 11344 1926.50
@@
   Retraction Reason (salient) 0.51 0.50 0 1 1
@@
   RetractionBot (ever during risk period) 0.59 0.49 0 1 1
   Number of Unique Editors (at risk start) 152.13 275.55 0 2343 34.50
@@
   Media Count (at risk start) 3.16 11.83 0 266 0
   Categories Count (at risk start) 3.43 3.92 0 56 2.00
   Reference Count (at risk start) 82.66 137.25 0 1416 36.00
Table 1: Summary statistics for the independent variables. RetractionBot is modeled as a time-dependent covariate in the extended Cox model; here we report whether a citation was ever monitored during its risk period (0/1).
To identify factors that are associated with faster or slower correction of retracted paper citations on Wikipedia, we examine three different groups of characteristics, as follows (see Table 1):
@@ 3.6.1 Paper Features
Paper characteristics are particularly important because they capture intrinsic properties of the publications that may be associated with differences in how quickly citations are corrected on Wikipedia. These data were retrieved from external scholarly databases, including OpenAlex Priem et al. (2022), Retraction Watch Oransky and Marcus (2012), and Altmetric Altmetric Support (2020).
@@
  • Open Access. Open access status is coded as a binary variable. This variable allows us to examine whether open-access availability is associated with correction timing on Wikipedia. Prior research suggests that open-access papers may receive greater visibility and more citations due to their unrestricted availability Piwowar et al. (2018); Teplitskiy et al. (2017). Zheng et al. (2024) found that open access may help speed up the detection and retraction of flawed articles in biochemistry Zheng et al. (2024).
@@
  • Academic Citation Count. Academic citations serve as a proxy for a paper’s scholarly impact and visibility. To avoid look-ahead bias, we measured this variable dynamically for each citation pair. Specifically, we only counted citations the paper had accumulated before the start of its risk period on Wikipedia, ensuring our measure of scholarly impact could not be influenced by the correction process itself. Using the OpenAlex API, we determined the number of citations a retracted paper had received up to the start of each citation’s unique risk period (i.e., the point after it was both retracted and cited on Wikipedia). While a paper’s higher citation count could increase its visibility and be associated with a faster correction on Wikipedia Van Noorden (2011); Wang and Barabási (2021), it could also create editorial inertia if its findings are widely accepted, resulting in a slower correction.
@@
  • Team Size. Team size is measured by counting the number of paper authors. Larger teams may contribute to greater research visibility and faster dissemination due to broader professional networks Li et al. (2022). However, prior studies suggest that increased team size can also complicate accountability, potentially leading to delays in retraction or correction processes Mongeon and Larivière (2016).
@@
  • Academic Domains. We include academic domains to examine whether the speed of citation correction varies across broad fields of research. Domains are derived from OpenAlex’s primary_topic field, which is based on Scopus’s ASJC classification Elsevier and assigned using OpenAlex’s topic classification methods OpenAlex (2024). We focus on four high-level domains: Health Sciences, Life Sciences, Physical Sciences, and Social Sciences. To incorporate these categorical variables into our regression model, we applied one-hot encoding and omitted Health Sciences as the reference category to avoid multicollinearity. Differences across domains may reflect disparities in editorial practices, public scrutiny, or disciplinary norms. For example, Health Sciences often receive greater media and public attention, which can accelerate the detection and correction of retracted work Furman et al. (2012). In contrast, corrections in the Social Sciences may lag due to longer publication cycles.
@@
  • Tweet Mention Count. Tweet mention count data are collected from Altmetric LLC Altmetric Support (2020). To ensure methodological consistency and avoid look-ahead bias, we measured the number of tweets up to the start of each citation’s unique “at risk” period. This count serves as a proxy for the paper’s online visibility and public exposure at the moment it became problematic on Wikipedia. While a higher count of these pre-existing tweets could signal a more prominent paper that editors are more likely to scrutinize, we acknowledge that the attention generated by the retraction event itself is another important mechanism. Although modeling this second effect is a valuable direction for future work, this study focuses on the role of a paper’s pre-existing visibility. Prior research highlights the role of online engagement in scientific controversies, finding that public attention on platforms like Twitter often spikes around the time a paper is retracted Peng et al. (2022). While news and blog mentions also serve as valid proxies for online engagement, we found that they were highly correlated with tweet mention counts. To avoid multicollinearity in our model, we therefore selected the Tweet mention count as the most representative measure.